Jetson Spacecraft Corp. shows the following information on its 2015 income statement: sales = $315973; costs = $211446; other expenses = $8759; depreciation expense = $19233; interest expense = $14143; taxes = $15748; dividends = $11614. In addition, you’re told that the firm issued $5750 in new equity during 2015 and redeemed $5210 in outstanding long-term debt. If net fixed assets increased by $21753 during the year, what was the addition to NWC?
Answer #1

Operating cash flow = EBIT + Depreciation - Taxes = (Sales- Costs-other expenses - depreciation) - taxes + depreciation
= (315973 -211446 -8759 -19233) -15748 + 19233 = 80020

Cash Flow from Assets = Cash flow to debtors + Cash flow to equity holders = (Interest - New Debt) + ( dividends- new Equity )
= (14143-5210) + (11614 - 5750 ) = 14797

Cash flow from Assets = OCF - Increase in Fixed = Increase in NWC
14,797 = 80020 - 21753 - Increase in NWC
Increase in NWC = 80020 -21753 -14,797 = 43,470

Please Discuss in case of Doubt

Best of Luck. God Bless
Please Rate Well